OpenStudy (tavia):

One card is selected at random from a deck of cards. Determine the probability that the card is a king. (Points : 4)

OpenStudy (anonymous):

4/52 = 2/26 = 1/13 Since there's 4 kings in a deck

OpenStudy (tavia):

ok thanks I thought there was a step I was supppose to do next but I have the same answer